OpenStudy (anonymous):

can someone explain the Newton-Raphson to me please??? i dont get it :S what does it work out and how??? thanks its for maths coursework so please help me *desperate*

OpenStudy (bahrom7893):

Well Newton-Raphson, also known as Newton's Method uses the slope of the function to approximate the value of the actual function.. I will have to draw this out and explain..

OpenStudy (bahrom7893):

http://www.twiddla.com/569411

OpenStudy (anonymous):

it would be great if you had an example to work with

OpenStudy (anonymous):

the idea is this: you want a zero of some function, say \[f(x)=x^3-4\] which is another way of saying you want the cube root of 4

OpenStudy (anonymous):

now make a guess

OpenStudy (anonymous):

use this x4+x-9=0 cause thats the equation i need to use do do the newton raphson method on

OpenStudy (anonymous):

ok fine we do that one

OpenStudy (anonymous):

\[f(x)=x^4+x-9\] make a guess

OpenStudy (anonymous):

it doesn''t even have to be a particularly good guess, just something that gets you relatively close to 0

OpenStudy (anonymous):

would you like me to make a guess?

OpenStudy (anonymous):

?? this is what ive done so far f(x)=x4+x-9 f’(x)=4x3+1 xr+1=xr - x4+x-9 4x3+1 the attachment is what ive got so far

OpenStudy (anonymous):

ok you have the idea. now just make a guess. i guess 1

OpenStudy (anonymous):

basically you look at an x value close to the solution , then t that x value, you go to the graph and draw a tangent line, and you use the x intercept of that tangent line as your new approximation ( which isnt always better )

OpenStudy (anonymous):

but guess what??? :S

OpenStudy (anonymous):

guess at some number that gives you something close to zero. i guess 1

OpenStudy (anonymous):

i did the tangent thing on autograph :)

OpenStudy (anonymous):

\[f(1)=-7\]

OpenStudy (anonymous):

a lousy guess but who cares

OpenStudy (anonymous):

\[f'(1)=5\]

OpenStudy (anonymous):

so my new guess is \[x_2=1-\frac{-7}{5}=1+\frac{7}{5}=2.2\]

OpenStudy (anonymous):

now i repeat the process with 2.2. in other words you make a guess at what you think the zero is

OpenStudy (anonymous):

call it g for guess. then you adjust by taking \[g-\frac{f(g)}{f'(g)}\] and that is your new guess.

OpenStudy (anonymous):

repeat until you think you are close enough, which you can tell because you will get the same decimals repeatedly

OpenStudy (anonymous):

its 2.4 its in the table of mydoc. but what does the newton raphson work out??? :S

OpenStudy (anonymous):

its repeating tangent lines, and it doesnt always work

OpenStudy (anonymous):

it gets you closer and closer to the zero. i can explain the geometry if you like

OpenStudy (anonymous):

imagine you have a point close to the zero you are looking for. say \[(x_1,y_1)\] where \[y_1\] is near zero, but not exactly 0 (otherwise \[x_1\] would be your answer.

OpenStudy (anonymous):

then find the equation of the line tangent to the graph. you do this by finding the slope, and using the point slope formula. you already have the point \[(x_1,y_1)\] and the slope is \[f'(x_1)\]

OpenStudy (anonymous):

the equation for the tangent line is \[y-y_1=f'(x_1)(x-x_1)\] now you want to know where this line crosses the x - axis (because you are trying to get to the zero of the function0 so set y = 0 and solve you get \[-y_1=f'(x)(x-x_1)\] \[x-x_1=-\frac{y_1}{f'(x_1)}\] \[x-x_1-\frac{f(x_1)}{f'(x_1)}\]

OpenStudy (anonymous):

im really sorry but i haven't got a clue what you are talking about :( i get it when your talking about tangents and a repeated root and getting closer to zero but i dont understand what the x1, y1) stuff?? im lost and what does the newton raphson actually work out??? roots???

OpenStudy (anonymous):

ok lets start at the very beginning and forget about tangents etc

OpenStudy (anonymous):

its just a linear approximation , dnt worry bout it too much

OpenStudy (anonymous):

what it works out is a numerical method to find the zeros of some function

OpenStudy (anonymous):

lets repeat your example because maybe it will make it clear. your equation is \[x^4+x-9=0\] which is the same as saying we want the zeros of \[f(x)=x^4+x-9\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok now here is the idea in a nutshell. it is a numerical method so you have to start somewhere. that is what i mean by "make a guess" guess some number \[x_1\] so that \[f(x_1)=0\] approximately

OpenStudy (anonymous):

so you approximate by a straight line ( which is easy to solve )

OpenStudy (anonymous):

it doesn't have to be a very good guess. i would guess \[x_1=1\] because i have no imagination and like integers

OpenStudy (anonymous):

with me so far?

OpenStudy (anonymous):

a linear eqn is alot easier to solve than a cubic

OpenStudy (anonymous):

as a matter of fact now that i look that was the first guess on the sheet you sent

OpenStudy (anonymous):

so far we have done nothing other than to make a guess at some number that gets us close to zero.

OpenStudy (anonymous):

doesnt matter what you first guess is , as long as its not a stationary point, or near one.

OpenStudy (anonymous):

unless the function has multiple turning points

OpenStudy (anonymous):

now the idea is this. see you close you are by finding \[f(x_1)\] in our case we get \[f(1)=1^4+1-9=-7\] so it was a rather lousy guess because -7 is not very close to zero. but whatever, we will adjust

OpenStudy (anonymous):

and this is how you adjust. find \[f'(x)=4x^3+1\] and now find \[f'(1)=5\]

OpenStudy (anonymous):

"adjust" :| lol

OpenStudy (anonymous):

so our first guess was 1 our adjusted guess will be \[1-\frac{f(1)}{f'(1)}\]

OpenStudy (anonymous):

in other words \[1-\frac{-7}{5}=2.2\] and that number is on your table in the sheet you have. that is the second guess

OpenStudy (anonymous):

well actually it is 2.4

OpenStudy (anonymous):

so now what? well now you do it again, this time with 2.4 instead of 1 \[2.4-\frac{f(2.4)}{f'(2.4)}\]

OpenStudy (anonymous):

now is where a calculator or spreadsheet would be handy because the numbers are annoying

OpenStudy (anonymous):

but the table on the left is the sequence of numbers you get by repeating the process.

OpenStudy (anonymous):

1 2.4 1.927895 1.700591 1.649107 1.646727 1.646722 1.646722

OpenStudy (anonymous):

and the numbers in the next column show that when you evaluate the function at these numbers you get closer and closer to 0

OpenStudy (anonymous):

-7 26.5776 6.74235427 1.064316798 0.045073275 9.23294E-05 3.89861E-10 0

OpenStudy (anonymous):

i am fairly sure that this did not answer your question. but i am not sure if the question is a mechanical one "how do we use this method?" or a conceptual one "how does it work""

OpenStudy (anonymous):

so what does the newton raphson work out? the roots? and if you use this method does the equation have to be ^4?

OpenStudy (anonymous):

copy paste that into the browser, it seemed to have broken the link :\

OpenStudy (anonymous):

Maybe a pic would help a bit.

OpenStudy (anonymous):

oh no it does not have to be 4th degree

OpenStudy (anonymous):

yes, it works out the roots

OpenStudy (anonymous):

you can use it to find square roots for example

OpenStudy (anonymous):

in fact it is more or less how your calculator finds them

OpenStudy (anonymous):

can be anything , e^(3x) + tan(x) +x^5 = 2

OpenStudy (anonymous):

if you want to find \[\sqrt{5}\] it is like saying find the roots of \[f(x)=x^2-5\] and you can use this method for that as well

OpenStudy (anonymous):

what elecengineer said. anything more or less

OpenStudy (anonymous):

@estudier nice picture!

OpenStudy (anonymous):

Thanks, did it help though?

OpenStudy (anonymous):

any mix of functions , inverse functions etc, they can all be linearly approximated. Thats the one thing you should understand/ appreciate , this method takes any messy equation and breaks it down to solving a linear equation

OpenStudy (anonymous):

i like the little balloon. problem is from the sheet this looks like a numerical methods class and not a calc one. there is precious little explanation on this paper

OpenStudy (anonymous):

i mean on the paper missmystery sent

OpenStudy (anonymous):

Yeah, methods, no background\theory, usual story...

OpenStudy (anonymous):

well yes, i now know how to work it out cause my friends did it for me on the doc and satellite you explained it very detailed :) thanks but i dunno what else to put for my coursework tho soz but my pc is really slow so it always takes a long while to reply :@

OpenStudy (anonymous):

Which bit of the coursework is still causing a problem?

OpenStudy (anonymous):

can you help me finish this sentence please??? But a disadvantage of using the decimal search is ..... and change in sign means..... thanks sorry pc is really really slow :(

OpenStudy (anonymous):

can you hepl me complete the two sentences pleasE??? change in signs means??? and a disadvantage of the decimal search please thanks sorry my pc is very slow at the moment and a while ago it didnt even let me go onto openstudy some message keeps poppin up :S

OpenStudy (anonymous):

A disadvantage of NR is that the sequence may not converge to a zero of the function. I'm not sure what u mean by change in signs (normally a sign change in relation to a gradient would mean that the curve had "turned").

OpenStudy (anonymous):

whats nr??? is that the same as decimal serach??? thanks and yes, i kinda remember now the change in sign is a turning point :)

OpenStudy (anonymous):

nr is Newton Raphson method, were you meaning the decimal search algorithm? Main problem with that is that the number of candidate solutions grows very fast and you need some kind of heuristics to keep things managable.

OpenStudy (anonymous):

thats for the decimal search right THANKS!!! :)